# On a conjecture of Peter Neumann on fixed points in permutation groups

### v1

@@ 7.5. Invariable generation of groups of Lie type
Assume now $S=\mathrm{P}\Omega_{2m+1}(r)$, so $m\geq 4$ and $r$ is odd. Note that $H$ is irreducible. (Indeed, both $x_{1}$ and $x_{2}$ fix a nondegenerate $1$-space, but one has square discriminant and the other has non-square discriminant.) Here [28, Corollary 3.4] does not apply and we use [27, Lemma 2.1 and Theorem 2.2] for the element $x_{1}$. By [27, Lemma 2.1] we have $\Phi^{*}_{2m}(r)>4m+1$; and since $m\geq 4$, by [27, Theorem 2.2] we deduce that $H=S$.
@@
Assume finally $S=\,^{2}E_{6}(r)$; we look at [14, Tables 3 and 10]. Note $|x_{1}|=\Phi^{*}_{18}(r)$ is equal to $19$ for $r=2$, and is at least $73$ for $r>2$ (by [27, Lemma 2.1]). The overgroups of $x_{1}$ are a subgroup of type $\mathrm{PSU}_{3}(r^{3})$, and $\mathrm{PSL}_{2}(19)$ (for $r=2$). None of these can contain $x_{2}$ and $H=S$. ∎
@@ 7.6. Generation of groups of Lie type by conjugates
In particular, it will be enough to show that the right-hand side of (7.4) is less than $1$, and in order to achieve this we will mostly use Lemma 2.1. Let also $\mathcal{M}=\mathcal{M}(x)$ be the set of maximal subgroups of $S$ containing $x$. Note that if $|\mathcal{M}(x)|=1$, say $\mathcal{M}(x)=\{M\}$, then the right-hand of (7.4) is at most $1/|S:M|<1$. When $S$ is classical, we will work with the cover of $S$ acting faithfully on the natural module, without changing notation.
@@
Assume now $S=\mathrm{SU}_{4}(r)$ and $|x|=\Phi^{*}_{2}(r^{2})$. We check the cases $r=2,3$ with GAP so assume $r\geq 4$. Then $|x|\geq 13$ by [27, Lemma 2.1]. We deduce from [7, p. 382] that the only overgroups of $x$ are the stabilizer $M_{1}$ of a totally singular $2$-space, the stabilizer $M_{2}$ of a decomposition into totally singular $2$-spaces, and subfield subgroups $\mathrm{Sp}_{4}(r).d$ and $\mathrm{SO}^{-}_{4}(r).e$, where d:=(2,r1)d:=(2,r-1) and $e:=(4,r+1)/2$ (there are $d$ and $e$ classes, respectively); the orthogonal subgroup being maximal only for $r$ odd. We have $\mathrm{fix}(x,S/M_{i})=1$ for $i=1,2$. Setting $M_{0}=\mathrm{Sp}_{4}(r)$, we can calculate $\mathrm{fix}(g,S/M_{3})^{2}/|S:M_{3}|$ precisely, similarly to (7.5); we see that this quantity is at most $1/r^{3}$, and the same holds for $M_{4}$. The same holds also for $M_{3}^{a}$ and $M_{4}^{a}$ with $a\in\mathrm{Aut}(S)$, therefore
@@
Assume now $S=\Omega^{+}_{2m}(r)$ with $m\geq 4$ even and $(m,r)\neq(4,2)$, so $|x|=\Phi^{*}_{2(m-1)}(r)$. Then $x$ stabilizes a unique nondegenerate $2$-space and $r+1$ nonsingular $1$-spaces. Moreover, $x$ lies in a unique extension field subgroup of type $\mathrm{GU}_{m}(r)$ if $m$ is even, and of type $\mathrm{O}_{m}(r)$ if $m$ is odd.
Let us first address the case $m=4$; see [7, pp. 402–403] for the list of maximal subgroups of $G$, taken from [36]. Letting $A=\langle\text{Inndiag}(\overline{S}),\tau\rangle$ where $\overline{S}=\mathrm{P}\Omega^{+}_{8}(r)$ and $\tau$ is a triality automorphism, we note that for $r$ odd (resp. $r$ even) there are six (resp. three) $\overline{S}$-classes of subgroups $\Omega_{7}(r)$ (resp. $\mathrm{Sp}_{6}(r)$) conjugate under $A$, four of which (resp. two of which) act (projectively) irreducibly via the spin module. (Note that $\tau$ does not lift to an automorphism of $S$; this is why we considered $\overline{S}$ here.) Let $M_{1}\leq S$ be the preimage of a representative for this $A$-class; we can take $M_{1}$ to be the stabilizer of some nonsingular $1$-space. Furthermore, there are three $\overline{S}$-classes of subgroups isomorphic to the stabilizer of a nondegenerate $2$-space of minus type, which are conjugate under $A$, and two of which are extension field subgroups of type $\mathrm{GU}_{4}(r)$. Let $M_{2}\leq S$ be the preimage of a representative for this $A$-class.
@@
7.7. Generation of alternating groups
@@ 8.1. Class $\mathcal{C}_{1}$
Subgroups $M$ in class $\mathcal{C}_{1}$ are stabilizers of certain subspaces $U$ of the natural module $V$. If $G$ preserves some nondegenerate form, then $U$ is nondegenerate, or totally singular, or a nonsingular $1$-space in orthogonal groups in even characteristic. What is more, if $U$ is nondegenerate, then $U$ is not similar to $U^{\perp}$. We also include the case where $G=\mathrm{SL}_{n}(q)$ and $M$ is the stabilizer of a flag or antiflag.
@@
Assume $G=\mathrm{SU}_{n}(q)$, so $M$ is the stabilizer of a nondegenerate subspace $U_{1}$ of dimension $m_{1}<n/2$, and set $m_{2}=n-m$. Write $V=U_{1}\perp U_{2}$ and, according to this decomposition, let $g=g_{1}\perp g_{2}$ where $g_{i}$ is as follows. If $m_{i}$ is odd, then $g_{i}$ acts irreducibly. If $m_{i}=2$ then $g_{i}$ is diagonal with distinct eigenvalues. If $m_{i}\neq 2$ is even, then $g_{i}$ acts irreducibly on each of a pair of complementary totally singular subspaces $A$ and $B$, and moreover $A\not\cong B$ as $\mathbb{F}_{q^{2}}\langle g_{i}\rangle$-modules. It is possible to choose $g_{1}$ and $g_{2}$ so that $\operatorname{det}(g)=1$, and such that $g_{1}$ and $g_{2}$ have no common eigenvalue, so $\mathrm{fix}(g,G/M)=1$.
@@
If $m_{i}$ is odd, then $g_{i}$ acts irreducibly on a nondegenerate hyperplane of minus type and trivially on a complement. If $m_{i}$ is even and $\delta_{i}=$-, then $g_{i}$ acts irreducibly. If $(m_{i},\delta_{i},q)=(4,+,2)$ then $g_{i}$ acts irreducibly on a nondegenerate $2$-space and trivially on a complement. If $m_{i}$ is even and $\delta_{i}=+$ and $(m_{i},q)\neq(4,2)$ then $g_{i}$ acts irreducibly on each of a pair of complementary totally singular subspaces $A$ and $B$, and moreover either $(m_{i},q)=(2,3)$ and $g_{i}=-1$, or $(m_{i},q)=(2,2)$ and $g_{i}=1$, or $A\not\cong B$ as $\mathbb{F}_{q}\langle g_{i}\rangle$-modules. It is possible to choose $g_{1}$ and $g_{2}$ so that $g\in\Omega_{n}^{\varepsilon}(q)$, and such that moreover $g_{1}$ and $g_{2}$ have no nontrivial irreducible submodule in common. In particular, we have $\mathrm{fix}(g,G/M)=1$ unless we are in one of the following cases:
@@
    $m_{1}$ and $m_{2}$ are odd, in which case $\mathrm{fix}(g,G/M)\leq q+1$ (the number of $1$-subspaces of a $2$-space).
@@
    $(m_{1},\delta_{1},q)=(2,+,3)$ and $m_{2}$ is odd, in which case $\mathrm{fix}(g,G/M)\leq 3$ (one plus the number of nondegenerate $1$-subspaces of $U_{1}$).
Assume now $M=\mathrm{Sp}_{n-2}(q)$ is the stabilizer of a nonsingular $1$-space with $q$ even. We take $g=g_{1}\perp g_{2}$ where $g_{1}$ acts irreducibly on a nondegenerate (n2)(n-2)-space and $g_{2}$ is trivial on a nondegenerate $2$-space $U$. Then $\mathrm{fix}(g,G/M)$ is the number of nonsingular $1$-spaces of $U$, so $\mathrm{fix}(g,G/M)\leq q+1$, and $(q+1)^{3}<q^{n-1}(q^{n}-1)\leq|G:M|$.
@@ 8.2. Class $\mathcal{C}_{2}$
Assume finally $n=4$. For $q=5$, let $g\in M$ be a regular (semisimple) element inducing a $3$-cycle on the spaces; concretely, for example, $g=(\lambda,1,1,-1)\tau$ where $\tau=(1,2,3)$ and $|\lambda|=4$. For $q\geq 7$, let $g\in\mathrm{GL}_{1}(q)^{4}\cap G$ be regular (semisimple). In both cases we have $|C_{G}(g)|<(q^{4}-1)/(q-1)$, and $(q^{4}-1)^{3}/(q-1)^{3}<|G:M|$.
@@
Assume now $G=\Omega^{+}_{n}(q)$ and $M$ of type $\mathrm{O}^{+}_{m}(q)\wr S_{t}$. In most cases, it will be sufficient to note that $M$ contains a suitable regular element. First note that a $t$-cycle $\tau$ of $S_{t}$ belongs to $\mathrm{SO}^{+}_{n}(q)\smallsetminus\Omega^{+}_{n}(q)$ if and only if $q\equiv 3\pmod{4}$ and $m\equiv 2\pmod{4}$ and $t$ is even, and otherwise $\tau\in\Omega^{+}_{n}(q)$. This follows from the fact that $\tau$ fixes complementary totally singular subspaces $A$ and $B$, and has determinant $(-1)^{m(t-1)/2}$ on both; recall that in this situation, $\Omega^{+}_{n}(q)$ consists of the elements whose determinant on $A$ (and $B$) is a square in $\mathbb{F}_{q}$ (see for example [39, Lemma 4.1.9]). Next, we have $(m,q)\neq(2,2),(2,3),(2,4),(4,2)$ by [7, Proposition 2.3.6]. In particular, if $(m,q)\neq(2,5),(4,3)$, we can choose $g=(g^{\prime},1,\ldots,1)\tau\in\Omega^{+}_{n}(q)$ where as an $\mathbb{F}_{q}\langle g^{\prime}\rangle$-module $V_{1}=A\oplus B$, where $A$ and $B$ are totally singular and irreducible, and $A\not\cong B$. Then $g$ has distinct eigenvalues, and the centralizer of $g$ in $\mathrm{O}^{+}_{n}(q)$ is isomorphic to the centralizer of a regular element of $\mathrm{GL}_{n/2}(q)$; therefore $|C_{G}(g)|<q^{n/2}$. Similarly to the $\mathrm{SL}_{n}(q)$ case, it can be easily checked that $q^{3n/2}<|G:M|$, which concludes the proof. We are left with $(m,q)=(2,5),(4,3)$. In these cases, setting $W=V_{1}\perp\cdots\perp V_{t-1}$, according to the decomposition $V=W\perp V_{t}$ we choose $g=g_{1}\perp g_{2}$, where $g_{1}=(g^{\prime},1,\ldots,1)\tau\in\mathrm{SO}^{+}_{n-m}(q)\smallsetminus\Omega^{+}_{n-m}(q)$ is defined as above (in particular, note that $g^{\prime}\in\mathrm{SO}^{+}_{m}(q)\smallsetminus\Omega^{+}_{m}(q)$ and $|g^{\prime}|=4$ for $(m,q)=(2,5)$, and $|g^{\prime}|=8$ for $(m,q)=(4,3)$), and $g_{2}$ is defined as follows: If $(m,q)=(2,5)$ then $g_{2}=g^{\prime}$, and if $(m,q)=(4,3)$ then as an $\mathbb{F}_{q}\langle g_{2}\rangle$-module $V_{t}=C\perp D$, where $C$ and $D$ are of minus type, $g_{2}$ is trivial on $D$ and has order $4$ on $C$. Next we divide into the two cases. Assume first $(m,q)=(4,3)$; note that $g$ is regular, and $|C_{G}(g)|<2q^{n/2-2}(q+1)^{2}$, which gives the conclusion for $n\geq 12$. For $n=8$, it is easy to see that $\mathrm{fix}(g,G/M)=1$ (since $V$ does not decompose as the sum of two equivalent $\mathbb{F}_{q}\langle g^{2}\rangle$-modules). Assume then $(m,q)=(2,5)$ and denote by $\lambda$ a generator of $\mathbb{F}_{5}^{\times}$. Then $g$ is not necessarily regular; specifically, $g$ is regular if and only if $t$ is odd. If $g$ is not regular, then for $\alpha=\lambda^{\pm 1}$ there are two $\alpha$-Jordan blocks, of sizes $1$ and the $5$-part of t1t-1. In particular, by working in $\mathrm{GL}_{n/2}(q)$ we see that $|C_{G}(g)|<q^{n/2+2}$. We have $q^{3n/2+6}<|G:M|$ and the result follows.

### v2

@@ 7.5. Invariable generation of groups of Lie type
Assume now $S=\mathrm{P}\Omega_{2m+1}(r)$, so $m\geq 4$ and $r$ is odd. Note that $H$ is irreducible. (Indeed, both $x_{1}$ and $x_{2}$ fix a nondegenerate 11-space, but one has square discriminant and the other has non-square discriminant.) Here [28, Corollary 3.4] does not apply and we use [27, Lemma 2.1 and Theorem 2.2] for the element $x_{1}$. By [27, Lemma 2.1] we have $\Phi^{*}_{2m}(r)>4m+1$; and since $m\geq 4$, by [27, Theorem 2.2] we deduce that $H=S$.
@@
Assume finally $S=\,^{2}E_{6}(r)$; we look at [14, Tables 3 and 10]. Note $|x_{1}|=\Phi^{*}_{18}(r)$ is equal to 1919 for $r=2$, and is at least $73$ for $r>2$ (by [27, Lemma 2.1]). The overgroups of $x_{1}$ are a subgroup of type $\mathrm{PSU}_{3}(r^{3})$, and $\mathrm{PSL}_{2}(19)$ (for $r=2$). None of these can contain $x_{2}$ and $H=S$. ∎
@@ 7.6. Generation of groups of Lie type by conjugates
In particular, it will be enough to show that the right-hand side of (7.4) is less than 11, and in order to achieve this we will mostly use Lemma 2.1. Let also $\mathcal{M}=\mathcal{M}(x)$ be the set of maximal subgroups of $S$ containing $x$. Note that if $|\mathcal{M}(x)|=1$, say $\mathcal{M}(x)=\{M\}$, then the right-hand of (7.4) is at most $1/|S:M|<1$. When $S$ is classical, we will work with the cover of $S$ acting faithfully on the natural module, without changing notation.
@@
Assume now $S=\mathrm{SU}_{4}(r)$ and $|x|=\Phi^{*}_{2}(r^{2})$. We check the cases $r=2,3$ with GAP so assume $r\geq 4$. Then $|x|\geq 13$ by [27, Lemma 2.1]. We deduce from [7, p. 382] that the only overgroups of $x$ are the stabilizer $M_{1}$ of a totally singular 22-space, the stabilizer $M_{2}$ of a decomposition into totally singular 22-spaces, and subfield subgroups $\mathrm{Sp}_{4}(r).d$ and $\mathrm{SO}^{-}_{4}(r).e$, where d:=(2,r1)d:=(2,r-1) and $e:=(4,r+1)/2$ (there are $d$ and $e$ classes, respectively); the orthogonal subgroup being maximal only for $r$ odd. We have $\mathrm{fix}(x,S/M_{i})=1$ for $i=1,2$. Setting $M_{0}=\mathrm{Sp}_{4}(r)$, we can calculate $\mathrm{fix}(g,S/M_{3})^{2}/|S:M_{3}|$ precisely, similarly to (7.5); we see that this quantity is at most $1/r^{3}$, and the same holds for $M_{4}$. The same holds also for $M_{3}^{a}$ and $M_{4}^{a}$ with $a\in\mathrm{Aut}(S)$, therefore
@@
Assume now $S=\Omega^{+}_{2m}(r)$ with $m\geq 4$ even and $(m,r)\neq(4,2)$, so $|x|=\Phi^{*}_{2(m-1)}(r)$. Then $x$ stabilizes a unique nondegenerate 22-space and $r+1$ nonsingular 11-spaces. Moreover, $x$ lies in a unique extension field subgroup of type $\mathrm{GU}_{m}(r)$ if $m$ is even, and of type $\mathrm{O}_{m}(r)$ if $m$ is odd.
Let us first address the case $m=4$; see [7, pp. 402–403] for the list of maximal subgroups of $G$, taken from [36]. Letting $A=\langle\text{Inndiag}(\overline{S}),\tau\rangle$ where $\overline{S}=\mathrm{P}\Omega^{+}_{8}(r)$ and $\tau$ is a triality automorphism, we note that for $r$ odd (resp. $r$ even) there are six (resp. three) $\overline{S}$-classes of subgroups $\Omega_{7}(r)$ (resp. $\mathrm{Sp}_{6}(r)$) conjugate under AA, four of which (resp. two of which) act (projectively) irreducibly via the spin module. (Note that $\tau$ does not lift to an automorphism of $S$; this is why we considered $\overline{S}$ here.) Let $M_{1}\leq S$ be the preimage of a representative for this AA-class; we can take $M_{1}$ to be the stabilizer of some nonsingular 11-space. Furthermore, there are three $\overline{S}$-classes of subgroups isomorphic to the stabilizer of a nondegenerate 22-space of minus type, which are conjugate under AA, and two of which are extension field subgroups of type $\mathrm{GU}_{4}(r)$. Let $M_{2}\leq S$ be the preimage of a representative for this AA-class.
@@
$7.$ Generation of alternating groups
@@ 8.1. Class $\mathcal{C}_{1}$
Subgroups $M$ in class $\mathcal{C}_{1}$ are stabilizers of certain subspaces $U$ of the natural module $V$. If $G$ preserves some nondegenerate form, then $U$ is nondegenerate, or totally singular, or a nonsingular 11-space in orthogonal groups in even characteristic. What is more, if $U$ is nondegenerate, then $U$ is not similar to $U^{\perp}$. We also include the case where $G=\mathrm{SL}_{n}(q)$ and $M$ is the stabilizer of a flag or antiflag.
@@
Assume $G=\mathrm{SU}_{n}(q)$, so $M$ is the stabilizer of a nondegenerate subspace $U_{1}$ of dimension $m_{1}<n/2$, and set $m_{2}=n-m$. Write $V=U_{1}\perp U_{2}$ and, according to this decomposition, let $g=g_{1}\perp g_{2}$ where $g_{i}$ is as follows. If $m_{i}$ is odd, then $g_{i}$ acts irreducibly. If $m_{i}=2$ then $g_{i}$ is diagonal with distinct eigenvalues. If $m_{i}\neq 2$ is even, then $g_{i}$ acts irreducibly on each of a pair of complementary totally singular subspaces AA and $B$, and moreover $A\not\cong B$ as $\mathbb{F}_{q^{2}}\langle g_{i}\rangle$-modules. It is possible to choose $g_{1}$ and $g_{2}$ so that $\operatorname{det}(g)=1$, and such that $g_{1}$ and $g_{2}$ have no common eigenvalue, so $\mathrm{fix}(g,G/M)=1$.
@@
If $m_{i}$ is odd, then $g_{i}$ acts irreducibly on a nondegenerate hyperplane of minus type and trivially on a complement. If $m_{i}$ is even and $\delta_{i}=$-, then $g_{i}$ acts irreducibly. If $(m_{i},\delta_{i},q)=(4,+,2)$ then $g_{i}$ acts irreducibly on a nondegenerate 22-space and trivially on a complement. If $m_{i}$ is even and $\delta_{i}=+$ and $(m_{i},q)\neq(4,2)$ then $g_{i}$ acts irreducibly on each of a pair of complementary totally singular subspaces AA and $B$, and moreover either $(m_{i},q)=(2,3)$ and $g_{i}=-1$, or $(m_{i},q)=(2,2)$ and $g_{i}=1$, or $A\not\cong B$ as $\mathbb{F}_{q}\langle g_{i}\rangle$-modules. It is possible to choose $g_{1}$ and $g_{2}$ so that $g\in\Omega_{n}^{\varepsilon}(q)$, and such that moreover $g_{1}$ and $g_{2}$ have no nontrivial irreducible submodule in common. In particular, we have $\mathrm{fix}(g,G/M)=1$ unless we are in one of the following cases:
@@
    $m_{1}$ and $m_{2}$ are odd, in which case $\mathrm{fix}(g,G/M)\leq q+1$ (the number of 11-subspaces of a 22-space).
@@
    $(m_{1},\delta_{1},q)=(2,+,3)$ and $m_{2}$ is odd, in which case $\mathrm{fix}(g,G/M)\leq 3$ (one plus the number of nondegenerate 11-subspaces of $U_{1}$).
Assume now $M=\mathrm{Sp}_{n-2}(q)$ is the stabilizer of a nonsingular 11-space with $q$ even. We take $g=g_{1}\perp g_{2}$ where $g_{1}$ acts irreducibly on a nondegenerate (n2)(n-2)-space and $g_{2}$ is trivial on a nondegenerate 22-space $U$. Then $\mathrm{fix}(g,G/M)$ is the number of nonsingular 11-spaces of $U$, so $\mathrm{fix}(g,G/M)\leq q+1$, and $(q+1)^{3}<q^{n-1}(q^{n}-1)\leq|G:M|$.
@@ 8.2. Class $\mathcal{C}_{2}$
Assume finally $n=4$. For $q=5$, let $g\in M$ be a regular (semisimple) element inducing a 33-cycle on the spaces; concretely, for example, $g=(\lambda,1,1,-1)\tau$ where $\tau=(1,2,3)$ and $|\lambda|=4$. For $q\geq 7$, let $g\in\mathrm{GL}_{1}(q)^{4}\cap G$ be regular (semisimple). In both cases we have $|C_{G}(g)|<(q^{4}-1)/(q-1)$, and $(q^{4}-1)^{3}/(q-1)^{3}<|G:M|$.
@@
Assume now $G=\Omega^{+}_{n}(q)$ and $M$ of type $\mathrm{O}^{+}_{m}(q)\wr S_{t}$. In most cases, it will be sufficient to note that $M$ contains a suitable regular element. First note that a $t$-cycle $\tau$ of $S_{t}$ belongs to $\mathrm{SO}^{+}_{n}(q)\smallsetminus\Omega^{+}_{n}(q)$ if and only if $q\equiv 3\pmod{4}$ and $m\equiv 2\pmod{4}$ and $t$ is even, and otherwise $\tau\in\Omega^{+}_{n}(q)$. This follows from the fact that $\tau$ fixes complementary totally singular subspaces AA and $B$, and has determinant $(-1)^{m(t-1)/2}$ on both; recall that in this situation, $\Omega^{+}_{n}(q)$ consists of the elements whose determinant on AA (and $B$) is a square in $\mathbb{F}_{q}$ (see for example [39, Lemma 4.1.9]). Next, we have $(m,q)\neq(2,2),(2,3),(2,4),(4,2)$ by [7, Proposition 2.3.6]. In particular, if $(m,q)\neq(2,5),(4,3)$, we can choose $g=(g^{\prime},1,\ldots,1)\tau\in\Omega^{+}_{n}(q)$ where as an $\mathbb{F}_{q}\langle g^{\prime}\rangle$-module $V_{1}=A\oplus B$, where AA and $B$ are totally singular and irreducible, and $A\not\cong B$. Then $g$ has distinct eigenvalues, and the centralizer of $g$ in $\mathrm{O}^{+}_{n}(q)$ is isomorphic to the centralizer of a regular element of $\mathrm{GL}_{n/2}(q)$; therefore $|C_{G}(g)|<q^{n/2}$. Similarly to the $\mathrm{SL}_{n}(q)$ case, it can be easily checked that $q^{3n/2}<|G:M|$, which concludes the proof. We are left with $(m,q)=(2,5),(4,3)$. In these cases, setting $W=V_{1}\perp\cdots\perp V_{t-1}$, according to the decomposition $V=W\perp V_{t}$ we choose $g=g_{1}\perp g_{2}$, where $g_{1}=(g^{\prime},1,\ldots,1)\tau\in\mathrm{SO}^{+}_{n-m}(q)\smallsetminus\Omega^{+}_{n-m}(q)$ is defined as above (in particular, note that $g^{\prime}\in\mathrm{SO}^{+}_{m}(q)\smallsetminus\Omega^{+}_{m}(q)$ and $|g^{\prime}|=4$ for $(m,q)=(2,5)$, and $|g^{\prime}|=8$ for $(m,q)=(4,3)$), and $g_{2}$ is defined as follows: If $(m,q)=(2,5)$ then $g_{2}=g^{\prime}$, and if $(m,q)=(4,3)$ then as an $\mathbb{F}_{q}\langle g_{2}\rangle$-module $V_{t}=C\perp D$, where $C$ and $D$ are of minus type, $g_{2}$ is trivial on $D$ and has order 44 on $C$. Next we divide into the two cases. Assume first $(m,q)=(4,3)$; note that $g$ is regular, and $|C_{G}(g)|<2q^{n/2-2}(q+1)^{2}$, which gives the conclusion for $n\geq 12$. For $n=8$, it is easy to see that $\mathrm{fix}(g,G/M)=1$ (since $V$ does not decompose as the sum of two equivalent $\mathbb{F}_{q}\langle g^{2}\rangle$-modules). Assume then $(m,q)=(2,5)$ and denote by $\lambda$ a generator of $\mathbb{F}_{5}^{\times}$. Then $g$ is not necessarily regular; specifically, $g$ is regular if and only if $t$ is odd. If $g$ is not regular, then for $\alpha=\lambda^{\pm 1}$ there are two $\alpha$-Jordan blocks, of sizes 11 and the 55-part of t1t-1. In particular, by working in $\mathrm{GL}_{n/2}(q)$ we see that $|C_{G}(g)|<q^{n/2+2}$. We have $q^{3n/2+6}<|G:M|$ and the result follows.